# FedDUAL: A Dual-Strategy with Adaptive Loss and Dynamic Aggregation for Mitigating Data Heterogeneity in Federated Learning

### v1

@@ Appendix D Convergence Analysis for Non-Convex Setting
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|\nabla F(w_{t})\|^{2}\leq\frac{C}{\sqrt{T}}$ (19)
where $C$ is a constant depending on the problem parameters.
We start with the same inequality given in  20.
$\mathbb{E}[F(w_{t+1})-F(w_{t})]\leq-\eta_{t}\|\nabla F(w_{t})\|^{2}+\frac{L\eta_{t}^{2}G^{2}}{2}$ (20)
Summing over $t=0$ to $T-1$:
$\sum_{t=0}^{T-1}\mathbb{E}[F(w_{t+1})-F(w_{t})]$ $\leq-\sum_{t=0}^{T-1}\eta_{t}\|\nabla F(w_{t})\|^{2}+\frac{LG^{2}}{2}\sum_{t=0}^{T-1}\eta_{t}^{2}$
$F(w^{*})-F(w_{0})$ $\leq-\sum_{t=0}^{T-1}\eta_{t}\|\nabla F(w_{t})\|^{2}+\frac{LG^{2}}{2}\sum_{t=0}^{T-1}\eta_{t}^{2}$
Using the learning rate $\eta_{t}=\frac{\eta}{\sqrt{T}}$ (constant for all $t$), we have:
$F(w^{*})-F(w_{0})$ $\leq-\frac{\eta}{\sqrt{T}}\sum_{t=0}^{T-1}\|\nabla F(w_{t})\|^{2}+\frac{LG^{2}\eta^{2}}{2}$
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(w_{t})\|^{2}$ $\leq\frac{\sqrt{T}}{\eta T}(F(w_{0})-F(w^{*}))+\frac{LG^{2}\eta}{2\sqrt{T}}$
Choosing $\eta=\sqrt{\frac{2(F(w_{0})-F(w^{*}))}{LG^{2}}}$ to balance the terms, we get:
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|\nabla F(w_{t})\|^{2}\leq\sqrt{\frac{2LG^{2}(F(w_{0})-F(w^{*}))}{T}}$
This is the required convergence rate for the non-convex setting of the proposed FedDUAL method.
Appendix E Discussion on convergence rates
E.1 Convex Setting
In the convex setting, The rate is optimal for first-order stochastic methods in convex optimization. The key points are:
1. The result is for the average iterate $\bar{w}_{T}$, which is common in convex stochastic optimization.
2. The rate depends on the initial distance to the optimum $\|w_{0}-w^{*}\|$, the initial suboptimality $F(w_{0})-F(w^{*})$, and the bound on the stochastic gradients $G^{2}$.
E.2 Non-Convex Setting
For the non-convex setting, The result is significantly different from the convex case:
1. We measure convergence in terms of the gradient norm, not the function value, as finding global minima is generally intractable in non-convex settings.
2. The result is for the average of the squared gradient norms, not the minimum.
3. This rate is also optimal for first-order stochastic methods in non-convex smooth optimization.
E.3 Comparison and Implications
1. Different Metrics: The convex and non-convex results use different metrics for convergence, which is standard in optimization theory. In the convex case, we bound the suboptimality of the function value, while in the non-convex case, we bound the average squared gradient norm.
2. Similarity in Rates: Both settings achieve an $O(1/\sqrt{T})$ rate, but the meaning is different. In the convex case, this rate applies to the optimality gap, while in the non-convex case, it applies to the average squared gradient norm.
3. Adaptive Loss and Wasserstein Barycenter: These novel components of our algorithm are captured in the constant factors and the bound $G^{2}$ on the stochastic gradients. They do not affect the asymptotic rate but improves the practical performance by reducing client heterogeneity.
4. Communication Efficiency: The $O(1/\sqrt{T})$ rate in both cases is in terms of the number of communication rounds, suggesting that our method is communication-efficient in both convex and non-convex settings.
5. Practical Implications: While the asymptotic rates are the same, the non-convex case may require more iterations to achieve a desired level of accuracy, as we’re working with squared gradient norms rather than function values.
In conclusion, our analysis differentiates between the convex and non-convex settings, providing appropriate convergence guarantees for each case. The proposed federated learning algorithm achieves optimal rates in both settings, demonstrating its theoretical efficiency and robustness to the convexity of the objective function. We achieve a convergence rate of $O\left(\frac{1}{\sqrt{T}}\right)$ towards a stationary point under smooth, non-convex conditions. This rate is on par with the traditional FedAvg method in non-convex settings (Yu et al., 2019).
@@ E.3 Comparison and Implications
Appendix F Ablation Study
F.1 Hyperparameter Analysis
In the proposed architecture, there are two key hyperparameters to consider: the scaling factor ($\gamma$) and the number of iterations used to compute the Wasserstein Barycenter. The detailed analysis of these hyperparameters is provided below.
The proposed FedDUAL approach utilizes dynamic server-side aggregation by applying the Wasserstein Barycenter concept to combine the weights of the final layers from local models. This iterative process involves a small positive constant ($\epsilon$) to determine the scaling factor ($\gamma$). To optimize performance, we conducted two sets of experiments on the FMNIST dataset with $\alpha=0.01$, each exploring a range of values for these crucial hyperparameters. The hyperparameter $\epsilon$ influences the sensitivity of the barycenter calculation to variations in Wasserstein distance. A smaller $\epsilon$ makes the barycenter more responsive to differences in Wasserstein distance, while a larger $\epsilon$ diminishes this sensitivity. This impacts how the barycenter integrates each distribution according to its distance from the current estimate. During the iterative update of the barycenter, $\epsilon$ affects the scaling factor $\gamma$ applied to each distribution. An excessively small $\epsilon$ can result in slow or potentially non-existent convergence due to minimal scaling factor, whereas a too-large $\epsilon$ may cause oversmoothing, reducing the barycenter’s effectiveness in accurately representing the distributions. For this setting we have fixed the number of iterations to compute Wasserstein Barycenter as 150. Figure 8 illustrates the test accuracy for different values of $\epsilon$, suggesting that larger $\epsilon$ values may degrade performance or even failed to achieve convergence. Figure 9 presents the corresponding learning curves for these settings.
The number of iterations in the Wasserstein Barycenter function is another critical hyperparameter that affects both the accuracy and efficiency of the barycenter computation. Generally, more iterations enhance convergence and accuracy, ensuring that the barycenter more closely approximates the optimal value. However, increasing the number of iterations also prolongs computation time, necessitating a balance between accuracy and efficiency. Finding the optimal number of iterations involves a trade-off: too few iterations may result in suboptimal outcomes, while too many can yield diminishing returns in accuracy. To achieve the best performance, begin with a reasonable default value, monitor convergence by observing changes in the barycenter, and adjust iteratively based on empirical results and available computational resources. For this setting, we fixed the $epsilon$ value as 0.0001, which yields the highest results in previous experiment. Figure 10 illustrates the test accuracy for different values of iterations to calculate Wasserstein Barycenter, suggesting that larger iterations may adversely affect performance. Figure 11 presents the corresponding learning curves for these settings.
From both experiments, we observe that the highest performance is achieved with $\epsilon=0.00001$ and 150 iterations. Therefore, to optimize performance, it is advisable to set $\epsilon$ to a smaller value while keeping the number of iterations between 100 and 150.
F.2 Experiment on Different Level of Data Heterogeneity
Figure 12 illustrates the accuracy of the proposed method and various baselines across different levels of data heterogeneity on the FMNIST dataset. In this context, heterogeneity is quantified by $\alpha$, with lower values indicating greater data heterogeneity. The results show that as $\alpha$ decreases, the test accuracy for all models declines. Remarkably, the proposed method consistently achieves the highest test accuracy and exhibits the slowest performance decline compared to other algorithms, demonstrating superior performance of the proposed method on varying degrees of non-IID data partitioning. The learning curve is presented in the Fig. 13.

### v2

@@ B.2 Discussion on Convergence Bounds
    $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\big[\|\nabla F(\theta_{t})\|^{2}\big]\leq\mathcal{O}\left(\frac{1}{\sqrt{T}}\right)+\epsilon_{\text{WB}}G,$
    consistent with the standard rate for non-convex optimization in federated learning. The additional $\epsilon_{\text{WB}}$ term quantifies the effect of Wasserstein Barycenter approximation, ensuring that its impact remains bounded.
Comparison with state of the art methods:
Standard algorithms such as FedAvg (McMahan et al., 2017) and FedProx (Li et al., 2020) achieve similar asymptotic rates ($\mathcal{O}(1/T)$ for convex and $\mathcal{O}(1/\sqrt{T})$ for non-convex), but suffer from significant degradation due to client drift under non-IID data. Specifically, FedAvg incurs heterogeneity-dependent terms of $\mathcal{O}(\zeta^{2}K^{2}/T)$ in non-convex settings, while FedProx improves this to $\mathcal{O}(\zeta^{2}K/T)$ through proximal regularization. Variance-reduced methods like SCAFFOLD (Karimireddy et al., 2020) achieve superior bounds with heterogeneity terms of only $\mathcal{O}(\zeta^{2}/T)$ (independent of local steps $K$) and variance scaling as $\mathcal{O}(1/nKT)$, but require maintaining control variates at each client. Our bounds show that FedDUAL matches these theoretical rates while introducing Wasserstein aggregation and adaptive loss weighting, providing robustness to heterogeneity without the additional memory and communication overhead of control variates where $T$ denotes the number of global communication rounds, $K$ is the number of local SGD steps performed at each client between communications, $n$ is the total number of participating clients, $\zeta$ is the client drift parameter measuring data heterogeneity, n is the total clients, and $\eta$ is the learning rate.
Appendix C Computational Complexity Analysis
We present a formal complexity analysis comparing the proposed method with baseline approaches for a single communication round involving $K$ clients, a model of dimension $d$, and $n$ denoting the number of parameters in the final layer.
Client-Side Complexity
FedAvg: The computational complexity per communication round is $\mathcal{O}(E\cdot B\cdot d)$, where $E$ denotes the number of local epochs, $B$ the batch size, and $d$ the model dimension. This cost primarily arises from the forward and backward passes performed during local training on each client.
FedProx: The per-round computational complexity is $\mathcal{O}(E\cdot B\cdot d+d)$, where the additional $\mathcal{O}(d)$ term accounts for computing the proximal regularization term $|\theta_{k}-\theta_{g}|^{2}$.
SCAFFOLD: The per-round complexity is $\mathcal{O}(E\cdot B\cdot d+2d)$, with the extra $\mathcal{O}(d)$ arising from storing and updating control variates used to correct client drift.
FedDUAL (Proposed): The per-round complexity is $\mathcal{O}(E\cdot B\cdot d+d)$, where the $\mathcal{O}(d)$ term corresponds to KL divergence computation comprising weight flattening, softmax operations, and divergence calculation, all linear in $d$. Thus, FedDUAL maintains the same asymptotic complexity as FedProx.
Please note that Our adaptive loss adds negligible client-side overhead compared to local training cost (as $E\cdot B\cdot d\gg d$).
Server-Side Complexity
FedAvg: The aggregation complexity per communication round is $\mathcal{O}(K\cdot d)$, corresponding to the simple averaging of model parameters across $K$ clients. FedProx/FedNova: The aggregation step also has complexity $\mathcal{O}(K\cdot d)$, as weight normalization and scaling operations remain linear in the model dimension.
SCAFFOLD: The aggregation complexity is $\mathcal{O}(K\cdot d)$, dominated by the averaging of client updates along with control variate terms.
FedDUAL (Proposed): The aggregation complexity is $\mathcal{O}(K\cdot d+I\cdot K\cdot n^{2})$, where $\mathcal{O}(K\cdot d)$ accounts for standard aggregation of lower-layer parameters and $\mathcal{O}(I\cdot K\cdot n^{2})$ arises from computing the Wasserstein barycenter for the final layers. Here, $I$ denotes the number of Sinkhorn iterations (typically 100–150), and $n$ represents the number of parameters in the last layers ($n\ll d$); for instance, in VGG-16, $d\approx 138$M total parameters, where the last two fully connected layers contain $n\approx 67.2$M parameters (two layers of $\times4096$ each), constituting approximately 48.7% of the total parameters.
Observation: While $n$ represents a substantial portion of the network, the Wasserstein aggregation complexity $\mathcal{O}(I\cdot K\cdot n^{2})$ remains tractable when applied selectively. For our setup:
  • VGG-16 on CIFAR-10: n=67,240,000n=67{,}240{,}000 (last two FC layers), $I=150$, $K=10$ (per round)
  • Wasserstein aggregation: $\sim 150\times 10\times(67{,}240{,}000)^{2}\approx 6.78\times 10^{17}$ operations
  • Standard aggregation: $10\times 138\times 10^{6}\approx 1.38\times 10^{9}$ operations
  • Overhead ratio: $\sim 4.9\times 10^{8}\times$ for the affected layers
However, this computational overhead is incurred only during the aggregation phase on the server, not during client-side training. Moreover, practical implementations use approximate Wasserstein distance computations (e.g., Sinkhorn iterations) which significantly reduce this theoretical complexity while maintaining the benefits of permutation-invariant aggregation.
subsubsection*Amortized Analysis Considering total time-to-convergence:
$T_{\text{total}}=R\times(T_{\text{client}}+T_{\text{server}})$ (86)
where $R$ is the number of rounds, $T_{\text{client}}$ is the per-round client training time, and $T_{\text{server}}$ is the per-round server aggregation time.
Since FedDUAL achieves convergence in $\sim$40% fewer rounds (refer to Fig. 4) while incurring $\sim$35% server-side overhead per round:
$\frac{T_{\text{FedDUAL}}}{T_{\text{FedAvg}}}=\frac{0.6R\times(T_{\text{client}}+1.35\times T_{\text{server}})}{R\times(T_{\text{client}}+T_{\text{server}})}=0.6\times\left(1+\frac{0.35\times T_{\text{server}}}{T_{\text{client}}+T_{\text{server}}}\right).$ (87)
In typical FL scenarios where client training dominates ($T_{\text{client}}\gg T_{\text{server}}$), we have $\frac{T_{\text{server}}}{T_{\text{client}}+T_{\text{server}}}\approx 0$, yielding:
$\frac{T_{\text{FedDUAL}}}{T_{\text{FedAvg}}}\approx 0.6\times(1+0)=0.6.$ (88)
Thus, the 35% server overhead becomes negligible, and FedDUAL achieves approximately 40% reduction in total wall-clock time due to faster convergence.
To empirically validate the assumption that $T_{client}\gg T_{server}$, we measured the average time required for a single client update and the total server update time per round. The results, summarized in Table 3, corroborate the assumption, demonstrating that client-side updates dominate the overall computational cost which further validate the general assumption made above.
Table 3: Wall clock time (in seconds) of the proposed method and the Fedavg algorithm for client and server side update.
CIFAR-10 FMNIST
T_client T_server T_client T_server
FedAvg 17.71 0.44 2.14 0.05
Proposed 27.49 1.94 2.62 1.95
Appendix D Runtime Analysis
Table 4 presents the wall-clock training time (in hours) of the proposed method compared to several widely used federated learning baselines on the CIFAR-10 and FMNIST datasets. This time is recorded for fixed number of rounds for all the algorithms (80 rounds for CIFAR10 dataset and 180 rounds for FMNIST dataset). As expected, the runtime varies across algorithms due to differences in their communication strategies, local computation overhead, and auxiliary regularization terms. Among the baselines, FedAvg achieves the lowest runtime, since it performs simple model averaging without any additional constraints or control variates. FedProx and FedBN incur a slight increase in runtime due to the introduction of proximal terms and batch normalization handling at the client side, respectively. FedNova and SCAFFOLD exhibit moderately higher runtimes, attributed to gradient normalization and control variate updates. MOON further increases computational cost because of its contrastive representation alignment loss, while shows additional overhead due to maintaining both personalized and shared components during optimization. FedDyn records the highest runtime among the existing methods, as its dynamic regularization term requires per-round model adjustment and additional global parameter updates. The proposed method, though computationally more expensive than the baselines, remains efficient considering the performance benefits (Higher accuracy and faster convergence) it achieves. Specifically, it requires 5.33 hours on CIFAR-10 and 3.01 holurs on FMNIST, which is competitive with other advanced methods such as FedDyn and FedPVR, while providing superior convergence and generalization performance. Overall, the proposed method achieves a favorable balance between computational cost and performance gain, validating its practicality in real-world federated learning scenarios.
We additionally report the wall-clock time (in hours) required to reach target accuracies of 0.40 for CIFAR-10 and 0.70 for FMNIST across all baselines and the proposed method, as summarized in Table 5. The results demonstrate that the proposed method consistently achieves the target accuracy in less time than the baselines, thereby confirming that the faster convergence observed in rounds (see Fig. 4) effectively translates into faster overall wall-clock convergence.
Table 4: Wall clock performance (in hours) of the proposed method and the baselines. * shows the respective method failed to converge.
CIFAR-10 FMNIST
FedAvg 2.66 1.75
FedProx 2.90 1.90
FedNova 3.10 *
FedBN * 1.80
FedDyn 4.89 2.91
MOON 3.80 2.40
SCAFFOLD * *
FedPVR 4.50 2.90
Proposed 5.33 3.01
Table 5: Wall-clock time to reach target accuracy (in hours). * indicates that the respective method failed to converge.
CIFAR-10 FMNIST
FedAvg 0.76 1.46
FedProx 1.61 1.90
FedNova 1.24 *
FedBN * 1.80
FedDyn 4.89 2.91
MOON 1.27 1.47
SCAFFOLD * *
FedPVR 1.31 1.21
Proposed 0.67 1.02
Appendix E Limitation and Future Work
While the proposed FedDUAL framework achieves superior performance and faster convergence under severe data heterogeneity, where several state-of-the-art methods such as FedNova, FedBN, and SCAFFOLD fail to converge, it introduces higher computational cost at the server due to the iterative calculation of the Wasserstein Barycenter. On the client side, FedDUAL remains lightweight and is also more communication-efficient than methods like SCAFFOLD and FedPVR, as it only requires transmitting model updates without additional control variates. Although this extra computational overhead is limited to the server, which typically has sufficient resources, future research will focus on reducing the computational burden of the Wasserstein Barycenter calculation by developing more efficient algorithms, aiming to maintain or even improve the performance of the proposed framework. Moreover, FedDUAL is inherently compatible with standard privacy-preserving techniques. In particular, its adaptive loss and dynamic aggregation can be seamlessly integrated with Differential Privacy by adding noise to client updates prior to aggregation, without modifying the core design. Investigating this integration with formal privacy guarantees remains an important direction for future research.
@@ Appendix E Limitation and Future Work
Figure 10: Learning curves of the proposed method and baselines with error bars.